OpenStudy (anonymous):

The sum of the measures of the interior angles of a polygon is five times the sum of the measures of its exterior angles, one angle at each vertex. How many sides does the polygon have?

OpenStudy (anonymous):

The sum of exterior angles of any polygon is 360 degrees. The sum of interior angles of any polygon is given by 180(n-2) degrees, n = number of sides. So the sum of the measure of interior angles, or 180(n-2) is equal to 5 times the sum of the measure of exterior angles, or 360. 180(n-2) = 5*360 Can you solve for n?

OpenStudy (anonymous):

i got n=12. is this correct?

OpenStudy (anonymous):

it is :)